# KAN4TSF: Are KAN and KAN-based models Effective for Time Series Forecasting?

### v1

@@ 5.3 Does KAN Outperform Linear?
Figure 3: Heatmap of the expert loads in the RMoK model on the Weather dataset. The value at location xy equal to 1 means that the gating network assigns the y-th variable to the x-th expert in all test samples.
5.4 Can KAN be Integrated into Other Methods?
In this section, we verify whether KAN can be integrated into the existing time series forecasting models as a plug-in to improve performance. We select iTransformer Liu et al. (2024b) as baseline and replace the linear projections of all the attention modules with different KAN variants and MoK layer. The experiments are conducted on the ETT datasets. For fair comparison, we set the same model hyperparameters in all experiments where hidden dim is 512 and layer number is 2, and search the best learning rate from 1e-2 to 1e-5 through a grid searching strategy, and repeat four times to report the average results in Table 4. While iTransformer with various KANs performs well only in ETTm1, iTransformer with MoK achieves the best performance in most cases on both ETTh1 and ETTm1. These experimental findings demonstrate that MoK is a successful form for integrating KAN into Transformer-based methods.
(a) Data visualization
(b) The weight of each input feature.
Figure 4: Visualization of KAN learning the periodicity in the data.
5.5 Are KAN-based Models Efficiency?
We report the size of model parameters, the training and infering speed of KAN-based methods and baselines on ETTh1 dataset with input length 96 and prediction length 720. We implement all methods in a unify code library with PyTorch Paszke et al. (2019), and the testing platform is a GPU server with NVIDIA A100 80GB GPUs. We set training batch size to 64, infer batch size to 1. The KAN, WKAN, TKAN, JAKN represent KANs with B-splines functions, Wavelet function (transform is mexican hat), Taylor polynomials (order=3), and Jacob polynomials (d is degree which sets to 4 or 6). As shown in Table 5, the running speed of KANs are affected by the specific implementation. KAN variant with Taylor polynomials can achieve close running efficiency to the Linear. With future hardware optimization, the efficiency of KANs will still be improved.
5.6 Are KAN-based Models Interpretable?
In this section, we aim to analyze the interpretability of RMoK for the time series forecasting task. First, we generate a heatmap to visualize the outputs of the gating network, which decide variable-expert assignment. This visualization demonstrates how the RMoK simplifies the multivariate time series forecasting task into multiple univariate forecasting subtasks. Then, we analyze what knowledge RMoK learns from real-world time-varying systems in each subtask.
@@ 5.6 Are KAN-based Models Interpretable?
Our proposed RMoK model uses a single MoK layer, which consists of gating network and KAN experts. The gating network outputs the matching score between variables and experts based on the input feature, and the top-$k$ matching experts are selected for each variable to predict the future state. We train an RMoK-S model with 4 experts on Weather dataset with 21 variables, then count the top-1 scores of all samples in the test set and generate a heatmap in Figure 3, where (x, y) = 1.0 means the y-th variable matches the x-th expert in all test samples. From this figure, although the matching score is related to the time-varying input data, the variable is still closely related to the specific expert. Therefore, we roughly approximate that the trained RMoK is a linear combination of several KANs, and RMoK treats the multivariate forecasting task as multiple univariate forecasting tasks. Although the above process is crude, this simplification helps us analyze the explainability of RMoK in complex real-world scenarios.
@@
After simplification, we try to analyze why KAN is effective in time series forecasting from the perspective of univariate and single-expert subtask. We use the temperature variable from Weather dataset which collects data every 10 minutes to train the KAN with B-splines function as forecasting model. We decompose and visualize data according to the trend, season, and resid terms through statsmodels  [^1] https://www.statsmodels.org/stable/index.html library. As shown in Figure 4, the temperature series has obvious daily periodicity, which is consistent with life experience. Then, we input the complete daily period (the past 144 time steps) data to predict the state of the next time step. We visualize the weight of each feature dimension of the trained RMoK. As shown in the Figure 4(b), there are three peaks in the feature weight, which are near 0, 72 and 144 time steps. 0 represents the temperature at the same time step of the previous day’s period, 144 represents the temperature at the adjacent moment, and 70 represents half a period. These three weight peaks correspond to the three zero points of the cosine function in one period. Finally, we conclude that RMoK can learn the periodicity of time series, which can preliminary explain its effectiveness in time series prediction.
@@ 6 Conclusion
This work discusses the effectiveness of KANs in time series forecasting. Due to various variants with different spline functions of KAN, we propose a single-layer mixture of KAN experts model (RMoK) to alleviate the distribution variation in time series. We experimentally compared KAN with existing baseline methods of various network architectures in terms of performance, integration, efficiency, and interpretability. Experimental results on seven real-world datasets show that RMoK performs in most metrics, it is sufficient to conclude that KAN and KAN-based models are effective in the time series forecasting task, and KAN can gain a place in the increasingly fierce model structure competition. The single-layer KAN-based model proposed in this work is only a preliminary attempt to introduce KAN into time series forecasting. We hope that our work can help future studies improve the performance and interpretability of KAN-based models.
Appendix A Implementation Details
A.1 Datasets
We conduct extensive experiments on seven real-world datasets, including ETT(h1, h2, m1, m2) Zhou et al. [2021], ECL, Traffic and Weather Lai et al. [2018], whose detail information is shown in Table 6. We follow the same data processing operations used in TimesNet Wu et al. [2023], where the training, validation, and testing sets are divided according to chronological order. The datasets we used cover multiple domains (electricity, transportation, and weather), which are sufficient to verify the generalizability of our method.
Table 6: The information of the seven time series forecasting datasets.
Dataset Variates Timesteps Granularity Domain
ETTh1,h2 7 17,420 1 hour Electricity
ETTm1,m2 7 69,680 15 min Electricity
ECL 321 26,304 1 hour Electricity
Traffic 862 17,544 1 hour Transportation
Weather 21 52,696 10 min Weather
A.2 Baselines
To comprehensively compare KAN with other baselines under different network structures, we select six well-known forecasting models as our baselines, including (1) Transformer-based methods: PatchTST and FEDformer; (2) CNN-based methods: TimesNet and SCINet; (3) Linear-based methods: RLinear and DLinear. We report the details of these baselines in Table 7.
Table 7: The information of baselines.
Name Type Cite
PatchTST Transformer Nie et al. [2023]
FEDformer Transformer Zhou et al. [2022]
TimesNet CNN Wu et al. [2023]
SCINet CNN Liu et al. [2022b]
RLinear Linear Li et al. [2023]
DLinear Linear Zeng et al. [2023]
A.3 Experimental Settings
We implement our method and baselines in a unify code library with PyTorch  [^2] https://github.com/pytorch/pytorch and Pytorch-lightning  [^3] https://github.com/Lightning-AI/pytorch-lightning (codes will be released after acceptance), and we conduct all experiments on a GPU server with NVIDIA A100 80GB GPUs. To prevent the sample drop in the test phase where the samples of the last batch are discarded, we set the test batch size to 1. Following previous works Zhou et al. [2021], Wu et al. [2023], we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of time series forecasting.
$\operatorname{MAE}(x,\hat{x})$ $=\frac{1}{N}\sum_{i=0}^{N}\left|x_{i}-\hat{x}_{i}\right|,$ (10)
$\operatorname{MSE}(x,\hat{x})$ $=\frac{1}{N}\sum_{i=0}^{N}\left(x_{i}-\hat{x}_{i}\right)^{2},$
where $x_{i}$ denotes the $i$-th ground truth, $\hat{x_{i}}$ represents the $i$-th predicted values, and $N$ represents the number of testing samples.
Appendix B Experimental Analysis
We report hyperparameter sensitivity and performance robustness experimental results to further evaluate the effectiveness of KAN-based modes in time series forecasting.
B.1 Hyperparameter Sensitivity
We conduct hyperparameter experiments to compare the performance between MoK and KAN layers. Results are shown in Table 8, where WavKAN, TaylorKAN and JacobKAN are KANs with wavelet functions, Taylor polynomials, and Jacobi polynomials. The type of spline function affects KAN’s ability to model time series, but there is currently a lack of theoretical guidance on how to choose the best suitable spline function, which can only be determined through experimental methods. Thus, MoK becomes a best practice for KAN, which does not require a lot of experiments to select the optimal spline function. This cocktail-like solution achieves better performance than a single KAN layer in most cases and achieves overall performance improvements on datasets with a large number of variables (such as Weather, Traffic, ECL).
Table 8: The performance of KAN-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
Setting with MoE w/o MoE
MoK-S MoK-B RWavKAN RTaylorKAN RJacobKAN
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.326 0.360 0.320 0.358 0.341 0.370 0.340 0.368 0.353 0.376
192 0.367 0.382 0.364 0.383 0.374 0.384 0.380 0.386 0.386 0.391
336 0.400 0.404 0.395 0.405 0.406 0.405 0.412 0.406 0.413 0.410
720 0.462 0.439 0.457 0.440 0.465 0.438 0.474 0.439 0.465 0.442
avg 0.389 0.396 0.384 0.397 0.397 0.399 0.402 0.400 0.404 0.405
ETTm2 96 0.176 0.261 0.176 0.261 0.179 0.265 0.181 0.264 0.177 0.263
192 0.244 0.305 0.240 0.302 0.247 0.309 0.247 0.306 0.246 0.308
336 0.306 0.346 0.299 0.342 0.311 0.349 0.308 0.345 0.309 0.348
720 0.405 0.404 0.397 0.401 0.412 0.408 0.405 0.401 0.406 0.403
avg 0.283 0.329 0.278 0.326 0.287 0.333 0.285 0.329 0.284 0.331
ETTh1 96 0.382 0.396 0.374 0.397 0.396 0.402 0.388 0.398 0.389 0.411
192 0.430 0.426 0.419 0.429 0.439 0.430 0.438 0.424 0.434 0.437
336 0.468 0.443 0.461 0.450 0.479 0.449 0.477 0.442 0.464 0.451
720 0.450 0.458 0.474 0.467 0.464 0.461 0.477 0.461 0.452 0.460
avg 0.433 0.431 0.432 0.436 0.444 0.435 0.445 0.431 0.435 0.440
ETTh2 96 0.313 0.357 0.301 0.353 0.321 0.358 0.297 0.344 0.310 0.352
192 0.397 0.409 0.379 0.405 0.407 0.413 0.387 0.401 0.396 0.405
336 0.441 0.446 0.432 0.446 0.452 0.450 0.434 0.441 0.437 0.441
720 0.453 0.461 0.446 0.463 0.487 0.477 0.448 0.458 0.445 0.456
avg 0.401 0.418 0.390 0.417 0.417 0.424 0.392 0.411 0.397 0.414
ECL 96 0.187 0.273 0.178 0.267 0.202 0.284 0.203 0.282 0.201 0.284
192 0.193 0.280 0.187 0.274 0.203 0.285 0.205 0.285 0.200 0.285
336 0.210 0.295 0.204 0.290 0.218 0.299 0.221 0.300 0.216 0.301
720 0.256 0.330 0.247 0.323 0.262 0.332 0.265 0.333 0.260 0.333
avg 0.211 0.294 0.204 0.288 0.221 0.300 0.224 0.300 0.220 0.301
Weather 96 0.175 0.225 0.171 0.221 0.180 0.228 0.182 0.232 0.178 0.230
192 0.224 0.267 0.220 0.263 0.227 0.268 0.235 0.274 0.229 0.271
336 0.284 0.308 0.277 0.302 0.284 0.308 0.291 0.312 0.288 0.310
720 0.366 0.358 0.360 0.354 0.368 0.360 0.370 0.360 0.368 0.359
avg 0.262 0.290 0.257 0.285 0.264 0.291 0.269 0.294 0.266 0.292

### v2

@@ 4.7 Case Study
Figure 4: Heatmap of the expert loads in the MoK on the Weather dataset. The value at location xy equal to 1 means that the gating network assigns the y-th variable to the x-th expert in all samples.
Figure 5: The weight of each input feature.
@@
In this section, we aim to analyze the interpretability of MMK for the time series forecasting task. First, we generate a heatmap to visualize the outputs of the gating network, which decide variable-expert assignment. The gating network outputs the matching score between variables and experts based on the input feature, and the top-$k$ matching experts are selected for each variable to predict the future state. We count the top-1 scores of all samples in the weather dataset and generate a heatmap in Figure 4, where (x, y) = 1.0 means the y-th variable matches the x-th expert in all samples. We can see that the variable is closely related to the specific expert, which indicates that MoK can effectively assign KANs to variables.
@@
Then, we use the temperature series in the weather dataset, which collects every 10 minutes to train the KAN model. We input one daily period (the past 144 time steps) of data to predict the state of the next time step. We visualize the weight of each feature dimension of the model. As shown in Figure 5, there are three peaks in the feature weight, which are near 0, 72, and 144 time steps. 0 represents the temperature at the same time step of the previous day’s period, 144 represents the temperature at the adjacent moment, and 70 represents half a period. We conclude that KAN can learn the periodicity of time series, which can preliminary explain its effectiveness in time series prediction.
@@ 5 Conclusion
The Kolmogorov-Arnold Network (KAN) seems to provide a new perspective for interpretable time series forecasting. In this paper, we evaluate the effectiveness of KANs in time-series forecasting tasks for the first time. First, we design the mixture-of-KAN layer, which uses a mixture-of-experts structure to assign variables to best-matched KAN experts. This module aggregates the advantages of various KAN’s variants. Then, we propose the multi-layer mixture-of-KAN network (MMK), which achieves excellent performance while retaining KAN’s ability to be transformed into a combination of symbolic functions. And we use some useful tricks to address the issues in the training stage. Finally, we compare MMK with many popular baselines on seven datasets. The effectiveness of KANs in multivariate time series forecasting has been proven by extensive experiments and visualizations. This work is only a preliminary attempt to introduce KAN into time series forecasting. We hope that our work can help future studies improve the performance and interpretability of KAN-based models.